# Get rich quick: exact solutions reveal how unbalanced initializations promote rapid feature learning

### v1

@@ B.1.3 Multi-Output
From our expression for $M(\beta)$ we will consider how it simplifies in settings of $\delta\to-\infty,0,\infty$ allowing us to gain intuition for $M_{\delta}(\beta)$.
@@
$M\to\begin{cases}\delta I&\delta\to-\infty.\\ \beta\beta^{\intercal}\otimes I+I\otimes\beta^{\intercal}\beta&\delta=0.\\ \delta I&\delta\to\infty.\end{cases}$ (102)
[…]
Computing $M^{-1}=\nabla^{2}q_{\delta}(\beta)$. Consistent with prior analyses, the natural next step of this derivation would be to compute the inverse of $M$ in order to find the potential $q_{\delta}$ indicative of the implicit bias. However, it is not straight forward to take the inverse of a Kronecker sum. Therefore, we turn to a simplification of our setting where we can precisely identify the influence $\delta$ has on the inductive bias .
Assumption B.4.
The singular value decomposition of the network function at initialization is $\mathrm{SVD}(\beta(0))=V\Lambda(0)U^{\intercal}$ where $\Lambda(0)$ is a diagonal matrix of strictly positive singular values and $U,V$ are the singular vectors of $X^{\intercal}Y$, ($\mathrm{SVD}(X^{\intercal}Y)=VSU^{\intercal}$),
Assumption B.5.
The input data $X^{\intercal}X$ is diagonalizable by $V$.
@@
Under Assumptions B.4 and B.5, if the gradient flow solution $\beta(\infty)$ satisfies $X^{\intercal}\beta(\infty)=Y$, then $\beta(\infty)=U\Lambda_{*}V^{\intercal}$ where $\Lambda_{*}$ solves the constrained optimization,
$\Lambda_{*}=\operatorname*{arg\,min}_{\begin{subarray}{c}\Lambda\in\mathrm{Diag}(\mathbb{R}^{d})\\ \Lambda_{ii}>0\quad\text{for }i=1,\ldots,i\end{subarray}}Q_{\delta}(\Lambda)\quad\mathrm{s.t.}\quad XU\Lambda V^{\intercal}=Y$ (103)
where $Q_{\delta}(\Lambda)=\sum_{i}^{h}q_{\delta}(\Lambda_{i})-\nabla q_{\delta}(\Lambda_{i}(0))\Lambda_{i}$ and
$q_{\delta}(x)=\frac{1}{4}\left(2x\sinh^{-1}\left(\frac{\sqrt{\eta_{a}\eta_{w}}2x}{\delta}\right)-\sqrt{\eta_{a}\eta_{w}4x^{2}+\delta^{2}}+\delta\right)$.
@@
Under assumptions B.4 and B.5, the network is said to be the task-aligned (Saxe et al. [10]), the eigenvectors of the network function $\beta$ are constant throughout training and equal to $U$ and $V$ respectivaly. We can therefore describe the dynamics of $\Lambda\in\mathrm{Diag}(\mathbb{R}^{d})$ with $\Lambda_{ii}>0\quad\text{for }i=1,\ldots,d$ where $\Lambda=V^{\intercal}\beta U$.
Rearranging the expressions for $W^{\intercal}W$ (equation 98) and $A^{\intercal}A$ (equation 99) in terms of the singular vectors of the task we get,
$A(t)^{\intercal}A(t)$ $=\frac{1}{\eta_{w}}U\left(\frac{\delta}{2}I+\sqrt{\eta_{a}\eta_{w}\Lambda^{2}+\left(\frac{\delta}{2}I\right)^{2}}\right)U^{\intercal}.$
$W^{\intercal}W$ $=\frac{1}{\eta_{a}}V\left(-\frac{\delta}{2}I+\sqrt{\eta_{a}\eta_{w}\Lambda^{2}+\left(\frac{\delta}{2}I\right)^{2}}\right)V^{\intercal}.$
Beginning with the network function dynamic $\beta$, rewrite the expression in terms of the eigenvectors $U$ and $V$
$\dot{\beta}$ $=-\left(\eta_{w}(X^{\intercal}X\beta-X^{\intercal}Y)A^{\intercal}A+\eta_{a}W^{\intercal}W(X^{\intercal}X\beta-X^{\intercal}Y)\right)$ (104)
$=-\left((X^{\intercal}XV\Lambda U^{\intercal}-VSU^{\intercal})U\Lambda_{AA}U^{\intercal}+V\Lambda_{WW}V^{\intercal}(X^{\intercal}XV\Lambda U^{\intercal}-VSU^{\intercal})\right)$ (105)
$=-\left(V((\Lambda-S)\Lambda_{AA}+\Lambda_{WW}(\Lambda-S))U^{\intercal}\right)$ (106)
Projecting $\dot{\beta}$ onto the singular vectors $V^{\intercal}$ and $U$, we now consider the dynamics of $\dot{\Lambda}$
${\dot{\Lambda}}$ $=\ -\left((\Lambda-S)\Lambda_{AA}+\Lambda_{WW}(\Lambda-S)\right)$ (107)
$=\ -\left((\Lambda-S)\left(\sqrt{\eta_{a}\eta_{w}\Lambda^{2}+\left(\frac{\delta}{2}I\right)^{2}}\right)+\left(\sqrt{\eta_{a}\eta_{w}\Lambda^{2}+\left(\frac{\delta}{2}I\right)^{2}}\right)(\Lambda-S)\right)$ (108)
Given that $\sqrt{\eta_{a}\eta_{w}\Lambda^{2}+\left(\frac{\delta}{2}I\right)^{2}}$ is a diagonal matrix and $(\Lambda-S)$ is also a diagonal matrix, the expression for $\dot{\eta_{a}\eta_{w}\Lambda}$ simplifies accordingly to
${\dot{\Lambda}}$ $=-M_{\Lambda}(\Lambda-S)$ (109)
with $M_{\Lambda}=\sqrt{\eta_{a}\eta_{w}4\Lambda^{2}+\delta^{2}I}.$
As layed out in Section A.4.1 the standard analysis for determining the implicit bias of a linear network through mirror flow, if the learning dynamics of the task aligned predictor $\Lambda$ can be expressed as a mirror flow for some strictly convex potential $\Phi_{\delta}(\beta)$,
$\dot{\Lambda}=-\left(\nabla^{2}\Phi_{\delta}(\Lambda)\right)^{-1}(\Lambda-S)$ (110)
Then, if the gradient flow solution $\beta(\infty)$ satisfies $X^{\intercal}\beta(\infty)=Y$, then $\beta(\infty)=U\Lambda_{*}V^{\intercal}$ where $\Lambda_{*}$ solves the constrained optimization
$\Lambda_{*}=\operatorname*{arg\,min}_{\begin{subarray}{c}\Lambda\in\mathrm{Diag}(\mathbb{R}^{d})\\ \Lambda_{ii}>0\quad\text{for }i=1,\ldots,h\end{subarray}}Q_{\delta}(\Lambda)\quad\mathrm{s.t.}\quad XU\Lambda V^{\intercal}=Y$ (111)
where $Q_{\delta}(\Lambda)=\Phi_{\delta}(\Lambda)-\nabla\Phi_{\delta}(\Lambda(0))^{\intercal}\Lambda$
The natural next step of this derivation is to compute the inverse of $M_{\Lambda}$ in order to find the potential $\Phi_{\delta}$ indicative of the implicit bias. $M_{\Lambda}=\sqrt{\eta_{a}\eta_{w}4\Lambda^{2}+\delta^{2}I}$ is a diagonal matrix therefore has a known inverse and
$(M_{\Lambda})^{-1}=\nabla^{2}\Phi_{\alpha}(\Lambda)=\left(\sqrt{\eta_{a}\eta_{w}4\Lambda^{2}+\delta^{2}I}\right)^{-1}$ (112)
Solving for the potential $\Phi_{\delta}(\Lambda)$. Consider the hypothesis that a scalar function $q(x)$ exists, allowing us to express $\Phi_{\delta}(\Lambda)$ = $\sum_{i=1}^{h}q_{\delta}(\Lambda_{i})+c_{\delta}$. Here, $c_{\delta}$ is a constant chosen to ensure that $\Phi_{\delta}(\Lambda_{ii})>0$ for all $\Lambda_{ii}>0$ for $i=1,\ldots,h$, and that $\Phi_{\delta}(0)=0$. The Hessian of this ansatz takes the form,
$\nabla^{2}\Phi_{\delta}(\Lambda)=\left(\sqrt{\eta_{a}\eta_{w}4\Lambda^{2}+\delta^{2}I}\right)^{-1}$ (113)
We now look for a function $q_{\delta}(x)$ and its derivatives,
$q_{\delta}(\Lambda_{i})=\frac{1}{4}\left(\sqrt{\eta_{a}\eta_{w}}2\Lambda_{i}\sinh^{-1}\left(\frac{\sqrt{\eta_{a}\eta_{w}}2\Lambda_{i}}{\delta}\right)-\sqrt{\eta_{a}\eta_{w}4\Lambda_{i}^{2}+\delta^{2}}+\delta\right)$ (114)
The hessian of $q_{\delta}(x)$ is
$\nabla^{2}\left(\sum_{i}^{h}q_{\delta}(\Lambda_{i})+c_{\delta}\right)=\left(\sqrt{\eta_{a}\eta_{w}4\Lambda^{2}+\delta^{2}}\right)^{-1}$ (115)
We find that the inductive bias is given by a hyperbolic entropy potential evaluated at the singular values of $\beta$. In the scenario of aligned networks, the distinction between $\ell^{1}$ and $\ell^{2}$ norms loses significance as the network consistently converges to the same solution. Nevertheless, the learning dynamics of $\beta$ will differ due to the potential $q_{\delta}$ smooth transition between an $\ell^{1}$ and $\ell^{2}$ penalty, distinguishing the rich and lazy regimes. This potential was initially identified as the inductive bias for diagonal linear networks by Woodworth et al. [14]. ∎
B.2 Multi-Layer
@@
We now consider the influence of depth by studying a depth-$(l+1)$ linear network, $f(x;\theta)=a^{\intercal}\prod_{i=1}^{l}W_{i}x$, where $W_{1}\in\mathbb{R}^{h\times d}$, $W_{i}\in\mathbb{R}^{h\times h}$ for $1<i\leq l$, and $a\in\mathbb{R}^{h}$. We assume that the dimensions $d=h$ and that all parameters share the same learning rate $\eta=1$. For this model the predictor coefficients are computed by the product $\beta=\prod_{i=1}^{l}W_{i}^{\intercal}a\in\mathbb{R}^{d}$. Similar to our analysis of a two-layer setting, we assume an isotropic initializations of the parameters.
@@
There exists a $\delta\in\mathbb{R}$ such that $aa^{\intercal}-W_{l}W_{l}^{\intercal}=\delta\mathbf{I}_{h}$ and for all $i\in[l-1]$ $W_{i+1}^{\intercal}W_{i+1}=W_{i}W_{i}^{\intercal}$.
This assumption can easily be achieved by setting $a=0$ and $W_{i}=\alpha O_{i}$ for all $i\in[l]$, where $O_{i}\in\mathbb{R}^{d\times d}$ is an random orthogonal matrix and $\alpha\geq 0$. In this case $\delta=-\alpha^{2}$. Further, notice this parameterization is naturally achieved in the high-dimensional limit as $d\to\infty$ under a standard Gaussian initialization with a variance inversely proportional with width. As in the two-layer setting, this structure of the initialization will remain conserved throughout gradient flow. We now show how two natural quantities of $\beta$, its squared norm $\|\beta\|^{2}$ and its outer product $\beta\beta^{\intercal}$, can always be expressed as polynomials of $\|a\|^{2}$ and $W_{1}^{\intercal}W_{1}$ respectively.
$\|\beta\|^{2}$ $=\|a\|^{2}\left(\|a\|^{2}-\delta\right)^{l}.$ (116)
@@
The norm of the regression coefficients is the product $\|\beta\|^{2}=a^{\intercal}\left(\prod_{i=1}^{l}W_{i}\right)\left(\prod_{i=1}^{l}W_{i}\right)^{\intercal}a$. Using the conservation of the initial conditions between consecutive weight matrices, $W_{i+1}^{\intercal}W_{i+1}=W_{i}W_{i}^{\intercal}$, we can express this telescoped product as $\|\beta\|^{2}=a^{\intercal}\left(W_{l}W_{l}^{\intercal}\right)^{d}a$. When plugging in the conservation between last two layers, this implies $\|\beta\|^{2}=a^{\intercal}\left(aa^{\intercal}-\delta I_{d}\right)^{d}a$, which expanded gives the desired result.
The outer product of the regression coefficients is $\beta\beta^{\intercal}=\left(\prod_{i=1}^{l}W_{i}\right)^{\intercal}aa^{\intercal}\left(\prod_{i=1}^{l}W_{i}\right)$. Using the conserved initial conditions of the last weights we can factor the outer product as the sum, $\beta\beta^{\intercal}=\left(\prod_{i=1}^{l}W_{i}\right)^{\intercal}W_{l}W_{l}^{\intercal}\left(\prod_{i=1}^{l}W_{i}\right)+\delta\left(\prod_{i=1}^{l}W_{i}\right)^{\intercal}\left(\prod_{i=1}^{l}W_{i}\right)$. Both these telescoping products factor using the conservation of the initial conditions between consecutive weight matrices giving the desired result. ∎
@@
$M=\left(W_{1}^{\intercal}W_{1}\right)^{l}+\|a\|^{2}\left(\sum_{i=0}^{l-1}(\|a\|^{2}-\delta)^{i}\left(W_{1}^{\intercal}W_{1}\right)^{l-1-i}\right).$ (118)
@@
Using a similar telescoping strategy used in the previous proofs we can derive the form of $M$, which we leave to the reader. ∎
@@
This theorem is a generalization of Proposition 1 derived in [9] for two-layer linear networks in the rich limit to deep linear networks in the rich limit. We find that the inductive bias, $Q(\beta)=(\tfrac{l+1}{l+2})\|\beta\|^{\frac{l+2}{l+1}}-\|\beta_{0}\|^{-\frac{l}{l+1}}\beta_{0}^{\intercal}\beta$, strikes a balance between attaining the minimum norm solution and preserving the initialization direction, which with increased depth emphasizes the latter.

### v2

@@ B.1.6 Simplifying $M$ through assumptions on $\Delta$
From this expression for $M(\beta)$ we can easily consider how it simplifies in limiting settings of $\delta$:
@@
$M\to\begin{cases}\delta\mathbf{I}_{dc}&\delta\to-\infty\\ \sqrt{\eta_{a}\eta_{w}\beta^{\intercal}\beta}\otimes\mathbf{I}_{d}+\mathbf{I}_{c}\otimes\sqrt{\eta_{a}\eta_{w}\beta\beta^{\intercal}}&\delta=0\\ \delta\mathbf{I}_{dc}&\delta\to\infty.\end{cases}$ (108)
As $\delta\to\pm\infty$, $M\to\delta\mathbf{I}_{dc}$, and the dynamics are lazy. In this limit, the dynamics of $\beta$ converge to the trajectory of linear regression trained by gradient flow and along this trajectory the NTK matrix remains constant. When $\delta=0$, $M=\sqrt{\eta_{a}\eta_{w}\beta^{\intercal}\beta}\otimes\mathbf{I}_{d}+\mathbf{I}_{c}\otimes\sqrt{\eta_{a}\eta_{w}\beta\beta^{\intercal}}$, and the dynamics are rich. Here the NTK changes in both magnitude and direction through training. In the next section we will attempt to better understand these dynamics for intermediate values of $\delta$ through the lens of a mirror flow.
B.1.7 Deriving a mirror flow for the singular values of $\beta$
For a matrix $\beta$, the dynamics of one of its singular values are given by $\dot{\sigma}=u^{\intercal}\dot{\beta}v$, where $u$ and $v$ are the corresponding left and right singular vectors. This equality can be derived from chain rule and the fact that $\|u\|=\|v\|=1$:
$\dot{\sigma}=\dot{u}^{\intercal}\beta v+u^{\intercal}\dot{\beta}v+u^{\intercal}\beta\dot{v}=\dot{u}^{\intercal}u\sigma+u^{\intercal}\dot{\beta}v+\sigma v^{\intercal}\dot{v}=u^{\intercal}\dot{\beta}v.$ (109)
In the last equality we used that fact that for any vector $z$ with a fixed norm, $\dot{\|z\|^{2}}=2\dot{z}^{\intercal}z=0$. Letting $\mathrm{diag}:\mathbb{R}^{d\times c}\to\mathbb{R}^{\min(d,c)}$ be the operator that, given a rectangular matrix, returns a vector of the elements on the main diagonal, we can then write,
$\dot{\lambda}=\mathrm{diag}(U^{\intercal}\dot{\beta}V)$ (110)
$\dot{\lambda}=-M\nabla_{\lambda}\mathcal{L}$ (111)
where $M$ is a diagonal matrix and $\nabla_{\lambda}\mathcal{L}$ is the gradient of the loss with respect to the singular values of $\beta$. Without loss of generality we consider $\eta_{a}=\eta_{w}=1$.
Lemma B.5.
Let $\Delta=\delta\mathbf{I}_{h}$. We then have that $\dot{\lambda}=-M\nabla_{\lambda}\mathcal{L}$, where $M\in\mathbb{R}^{\min(d,c)\times\min(d,c)}$ is a diagonal matrix with
$M_{ii}=\begin{cases}\sqrt{\delta^{2}+4\lambda_{i}^{2}}&i\leq\min(d,h,c)\\ 0&\mathrm{otherwise}\end{cases}$ (112)
Proof.
First note that
$\dot{\lambda}$ $=\mathrm{diag}(U^{\intercal}\dot{\beta}V)$ (113)
$=-\mathrm{diag}\left(U^{\intercal}\left[X^{\intercal}(X\beta-Y)A^{\intercal}A+W^{\intercal}WX^{\intercal}(X\beta-Y)\right]V\right)$ (114)
$=-\mathrm{diag}\left(U^{\intercal}X^{\intercal}(X\beta-Y)V\Sigma_{A}^{2}+\Sigma_{W}^{2}U^{\intercal}X^{\intercal}(X\beta-Y)V\right)$ (115)
where we let $W^{\intercal}W=U\Sigma_{W}^{2}U^{\intercal}$ and $A^{\intercal}A=V\Sigma_{A}^{2}V^{\intercal}$, using the fact that, under $\Delta=\mathbf{I}_{h}$, the eigenvectors of $A^{\intercal}A$ are the right singular vectors of $\beta$ and the eigenvectors of $W^{\intercal}W$ are the left singular vectors of $\beta$. This expression rewrites as
$\dot{\lambda}=-M\mathrm{diag}\left(U^{\intercal}X^{\intercal}(X\beta-Y)V\right)$ (116)
where $M\in\mathbb{R}^{\min(d,c)\times\min(d,c)}$ is a diagonal matrix with $M_{ii}=(\Sigma_{A}^{2})_{ii}+(\Sigma_{W}^{2})_{ii}$. For $i\leq\min(d,h,c)$, one can show that $M_{ii}=\sqrt{\delta^{2}+4\lambda_{i}^{2}}$. This is because for $i\leq\min(d,h,c)$, $(\Sigma_{A}^{2})_{ii}=(\Sigma_{W}^{2})_{ii}+\delta$ from the conservation law and $(\Sigma_{W}^{2})_{ii}(\Sigma_{A}^{2})_{ii}=\lambda_{i}^{2}$ from the definition of $\lambda$. Together this implies $(\Sigma_{W}^{2})_{ii}\left(\delta+(\Sigma_{W}^{2})_{ii}\right)=\lambda_{i}^{2}$, which is a quadratic equation in $(\Sigma_{W}^{2})_{ii}$. If $h<\min(d,c)$ then $M_{ii}=0$ for $i>\min(d,c)$ accounting for rank deficiency of both $A$ and $W$ in this case. Additionally, in our setting of MSE loss, it is straightforward to show that
$\frac{\partial\mathcal{L}}{\partial\lambda_{i}}=\left(U^{\intercal}X^{\intercal}(X\beta-Y)V\right)_{ii}$ (117)
We then have that $\nabla_{\lambda}\mathcal{L}=\mathrm{diag}\left(U^{\intercal}X^{\intercal}(X\beta-Y)V\right)$, which, combined with our expression for $M$, completes the proof. ∎
Leveraging Lemma B.5, we can show that the singular values of $\beta$ evolve under a mirror flow in the following theorem.
@@ B.1.7 Deriving a mirror flow for the singular values of $\beta$
Let $\Delta=\delta\mathbf{I}_{h}$ and assume $h\geq\min(d,c)$ and $\delta\neq 0$. We then have that the dynamics of $\lambda$, the singular values of $\beta$, are given by the mirror flow
$\dot{\lambda}=-\left(\nabla^{2}\Phi_{\delta}(\lambda)\right)^{-1}\nabla_{\lambda}\mathcal{L},$ (118)
where $\Phi_{\delta}(\lambda)=\sum_{i=1}^{\min(d,c)}q_{\delta}(\lambda_{i})$ and $q_{\delta}$ is the hyperbolic entropy potential
$q_{\delta}(x)=\frac{1}{4}\left(2x\sinh^{-1}\left(\frac{2x}{|\delta|}\right)-\sqrt{4x^{2}+\delta^{2}}+|\delta|\right).$ (119)
@@
When $\Delta=\delta\mathbf{I}_{h}$, then by Lemma B.5 the dynamics of the singular values of $\beta$ can be expressed as $\dot{\lambda}=-M\nabla_{\lambda}\mathcal{L}$. Furthermore, when $h\geq\min(d,c)$ and $\delta\neq 0$, we have that $M=\sqrt{\delta^{2}+4\lambda^{2}}\mathbf{I}_{\min(d,c)}$, where $\lambda^{2}$ is element-wise, which is always invertible. Observe, this expression for $M$ is the inverse Hessian of the potential $\Phi_{\delta}(\lambda)=\sum_{i}q_{\delta}(\lambda_{i})$ for $q_{\delta}$ specified in the theorem statement. Thus, the dynamics for the singular values are the mirror flow $\dot{\lambda}=-\left(\nabla^{2}\Phi_{\delta}(\lambda)\right)^{-1}\nabla_{\lambda}\mathcal{L}$. ∎
Theorem B.6 implies that the dynamics for the singular values of $\beta$ can be described as a mirror flow with a $\delta$-dependent potential. This potential was first identified as the inductive bias for diagonal linear networks by Woodworth et al. [14]. Termed hyperbolic entropy, this potential smoothly interpolates between an $\ell^{1}$ and $\ell^{2}$ penalty on the singular values for the rich ($\delta\to 0$) and lazy ($\delta\to\pm\infty$) regimes respectively. Unfortunately, in our setting we cannot adapt our mirror flow interpretation into a statement on the inductive bias at interpolation because the singular vectors evolve through training. If we introduce additional assumptions — specifically, whitened input data ($X^{\intercal}X=\mathbf{I}_{d}$) and a task-aligned initialization such that the singular vectors of $\beta_{0}$ are aligned with those of $\beta_{*}$ — we can ensure that the singular vectors remain constant and thus derive an inductive bias on the singular values. However, in this setting the dynamics decouple completely, implying there is no difference between applying an $\ell^{1}$ or $\ell^{2}$ penalty on the singular values. Consequently, even though the dynamics will depend on $\delta$, the final interpolating solution will be independent of $\delta$, making a statement on the inductive bias insignificant.
B.2 Deep linear networks
@@
We now consider the influence of depth by studying a depth-$(L+1)$ linear network, $f(x;\theta)=a^{\intercal}\prod_{l=1}^{L}W_{l}x$, where $W_{1}\in\mathbb{R}^{h\times d}$, $W_{l}\in\mathbb{R}^{h\times h}$ for $1<l\leq L$, and $a\in\mathbb{R}^{h}$. We assume that the dimensions $d=h$ and that all parameters share the same learning rate $\eta=1$. For this model the predictor coefficients are computed by the product $\beta=\prod_{l=1}^{L}W_{l}^{\intercal}a\in\mathbb{R}^{d}$. Similar to our analysis of a two-layer setting, we assume an isotropic initializations of the parameters.
@@
There exists a $\delta\in\mathbb{R}$ such that $aa^{\intercal}-W_{L}W_{L}^{\intercal}=\delta\mathbf{I}_{h}$ and for all $l\in[L-1]$ $W_{l+1}^{\intercal}W_{l+1}=W_{l}W_{l}^{\intercal}$.
This assumption can easily be achieved by setting $a=0$ and $W_{l}=\alpha O_{l}$ for all $l\in[L]$, where $O_{l}\in\mathbb{R}^{d\times d}$ is an random orthogonal matrix and $\alpha\geq 0$. In this case $\delta=-\alpha^{2}$. Further, notice this parameterization is naturally achieved in the high-dimensional limit as $d\to\infty$ under a standard Gaussian initialization with a variance inversely proportional with width. As in the two-layer setting, this structure of the initialization will remain conserved throughout gradient flow. We now show how two natural quantities of $\beta$, its squared norm $\|\beta\|^{2}$ and its outer product $\beta\beta^{\intercal}$, can always be expressed as polynomials of $\|a\|^{2}$ and $W_{1}^{\intercal}W_{1}$ respectively.
Lemma B.8.
For a depth-$(L+1)$ linear network with square width ($d=h$) and isotropic initialization, then for all $t\geq 0$,
$\|\beta\|^{2}$ $=\|a\|^{2}\left(\|a\|^{2}-\delta\right)^{L},$ (120)
@@
The norm of the regression coefficients is the product $\|\beta\|^{2}=a^{\intercal}\left(\prod_{l=1}^{L}W_{l}\right)\left(\prod_{l=1}^{L}W_{l}\right)^{\intercal}a$. Using the conservation of the initial conditions between consecutive weight matrices, $W_{l+1}^{\intercal}W_{l+1}=W_{l}W_{l}^{\intercal}$, we can express this telescoped product as $\|\beta\|^{2}=a^{\intercal}\left(W_{L}W_{L}^{\intercal}\right)^{d}a$. When plugging in the conservation between last two layers, this implies $\|\beta\|^{2}=a^{\intercal}\left(aa^{\intercal}-\delta\mathbf{I}_{h}\right)^{d}a$, which expanded gives the desired result.
The outer product of the regression coefficients is $\beta\beta^{\intercal}=\left(\prod_{l=1}^{L}W_{l}\right)^{\intercal}aa^{\intercal}\left(\prod_{l=1}^{L}W_{l}\right)$. Using the conserved initial conditions of the last weights we can factor the outer product as the sum, $\beta\beta^{\intercal}=\left(\prod_{l=1}^{L}W_{l}\right)^{\intercal}W_{L}W_{L}^{\intercal}\left(\prod_{l=1}^{L}W_{l}\right)+\delta\left(\prod_{l=1}^{L}W_{l}\right)^{\intercal}\left(\prod_{l=1}^{L}W_{l}\right)$. Both these telescoping products factor using the conservation of the initial conditions between consecutive weight matrices giving the desired result. ∎
@@
$M=\left(W_{1}^{\intercal}W_{1}\right)^{L}+\|a\|^{2}\left(\sum_{l=0}^{L-1}(\|a\|^{2}-\delta)^{l}\left(W_{1}^{\intercal}W_{1}\right)^{L-1-l}\right).$ (122)
@@
Using a similar telescoping strategy used in the previous proof we obtain the form of $M$. ∎
@@
This theorem is a generalization of Proposition 1 derived in [9] for two-layer linear networks in the rich limit to deep linear networks in the rich limit. We find that the inductive bias, $Q(\beta)=(\tfrac{L+1}{L+2})\|\beta\|^{\frac{L+2}{L+1}}-\|\beta_{0}\|^{-\frac{L}{L+1}}\beta_{0}^{\intercal}\beta$, strikes a depth-dependent balance between attaining the minimum norm solution and preserving the initialization direction.